OpenStudy (anonymous):

The circumference of a circle with a diameter of 3.73 in is _____in. Round answer to nearest hundredth.

OpenStudy (swissgirl):

Ok the circumference of a circle is \( 2 \pi r\) And the diameter= 2*r

OpenStudy (anonymous):

I got the answer but I need to know exactly what it is moved 2 decimal places

OpenStudy (swissgirl):

so what did you get?

OpenStudy (anonymous):

23.4362811959

OpenStudy (swissgirl):

Okkk so basically The diameter is 2*r so d=2r The circumference of a circle is \( 2 \pi r = \pi *d\) d= 3.73 so the circumference= 3.73 * pi= 11.7181406 which we round to 11.72

OpenStudy (anonymous):

so is that the answer

OpenStudy (swissgirl):

Should be :)

OpenStudy (anonymous):

thanks

OpenStudy (swissgirl):

You are welcome :)